OpenStudy (anonymous):

Given the equation y − 4 = three fourths(x + 8) in point-slope form, identify the equation of the same line in standard form. −three fourthsx + y = 10 3x − 4y = −40 y = three fourthsx + 12 y = three fourthsx + 10

OpenStudy (tkhunny):

Isn't "Standard Form" like this: (Integer)x + (Integer)y = (integer)? There's only one choice taht looks like that.

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

Well, then we need to get rid fo the fractions, right? y − 4 = ¾(x + 8) Multiply by 4 4y − 16 = 3(x + 8) What's next?